OpenStudy (anonymous):

Solve –2(2x – 3) = –6

OpenStudy (anonymous):

remove parentheses using distributive law

jhonyy9 (jhonyy9):

-4x=-12 x=3

OpenStudy (anonymous):

-2(2x-3)=-6 -4x+6=-6 6+6=4x 12=4x 3=x

OpenStudy (anonymous):

\[-2\times 2x+(-2)\times (-3)=-6\] \[-4x+6=-6\] subtract 6 from both sides \[-4x+6-6=-6-6\] \[-4x=-12\] divide both sides by -4 \[\frac{-4x}{-4}=\frac{-12}{-4}\] \[x=3\]